# Gravitational wave probes of particle dark matter: a review

### v4

@@ II.2 Axions
Axions were originally proposed as a solution to the strong CP problem in QCD Peccei and Quinn (1977a, b), and more generally, axion-like particles arise in many extensions of the standard model. If sufficiently light and weakly interacting, they can also constitute a viable component of DM. A key feature of axions is their ability to couple to standard-model fields, in particular photons, through a Chern–Simons interaction. This coupling is described by the Lagrangian Chern and Simons (1974):
@@
the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived Nagano et al. (2019):
@@
Eq. 11 implies that in the presence of an axion field, the phase velocities of circularly polarized photons in interferometers are modulated. In order to detect this effect in LIGO, Virgo and KAGRA instruments, additional but simplistic optical components are required in order to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-polarizations) Nagano et al. (2021). In practice, GW interferometers use linearly polarized light ($p$-polarized), which becomes partially $s$-polarized due to axion-induced modulation. Similar deviations in polarization can also be probed with pulsar timing arrays Xue et al. (2024); Porayko et al. (2025), and experimental constraints on axions from them will be discussed in Section III.3.5. Notably, ground-based interferometers and pulsar timing arrays are sensitive to very different axion mass ranges, roughly $10^{-13},10^{-11}]$ eV and $10^{-23},10^{-21}]$ eV, respectively, but can achieve comparable absolute constraints on the axion-photon coupling $g_{\text{a}\gamma}$, which vary strongly with the axion mass. However, no constraints exist yet for this effect from ground-based interferometers.
@@ II.3 Dilaton scalar dark matter
Another class of ultralight dark matter candidates are dilatons, scalar fields that can couple to standard-model particles and induce time-dependent variations in fundamental constants Preskill et al. (1983); Abbott and Sikivie (1983); Dine and Fischler (1983); Cho and Keum (1998); Cho and Kim (2009); Arvanitaki et al. (2015b); Stadnik and Flambaum (2015a, b, 2016). In particular, a dilaton field can cause the electron mass and other physical constants to oscillate at the dilaton frequency. Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki et al. (2015b); Derevianko (2018); Vermeulen et al. (2021):
@@
where $\vec{x}$ is a position vector, $\vec{k}=m_{\rm DM}\vec{v}_{0}$ is the wave vector, and $\Psi$ is a phase factor, and has a corresponding Lagrangian $\mathcal{L}_{\mathrm{int}}$ Vermeulen et al. (2021):
@@ II.3.1 Changes to size and indices of refraction of solids
The oscillations in fundamental constants induced by a dilaton field lead to modulations of the Bohr radii of atoms in interferometer components Grote and Stadnik (2019), leading to measurable changes in the size and refractive index of the beam splitter Vermeulen et al. (2021), and the arm mirrors Göttel et al. (2024). Oscillations in the beam splitter size cause photons traveling down each arm to traverse slightly different paths along its surface, producing a differential phase shift. Initially, it was assumed that the beam-splitter effect was the dominant contribution Vermeulen et al. (2021), but more recent work has shown that small differences in the thicknesses of the mirrors in each arm also lead to non-negligible path-length variations in the interferometer Göttel et al. (2024).
@@
In the context of GW interferometers, this effect translates into a differential displacement of the beam splitter along the two arms. For the $x$-arm, assuming light incidence at 45 degrees and 50% mirror reflectivity, the induced change in arm length, $\delta L_{x}$, can be computed as in Grote and Stadnik (2019):
@@
where $w$ is the thickness of the mirrors, $n$ is the index of the beamsplitter, and $l$ is the length of the beamsplitter. Along the $y$-arm, the change in arm length is Vermeulen et al. (2021):
@@
We have only considered the contribution arising from the change in length of the beamsplitter, not from the change in index of refraction, because the former is much larger than the latter Grote and Stadnik (2019).
In addition to dilatons interacting with atoms in the beamsplitter, the reference cavity of the detector can provide a region in which dilatons interact appreciably with light, allowing for potential detection. Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko (2018); Geraci et al. (2019); Hall and Aggarwal (2022). The strain amplitude of such a signal, which would be distinct from that in Eq. 26, can be written as:
@@
The effects described in Eq. 26 and Eq. 27 can both produce measurable signals in interferometers. Currently, however, only the spatial strain affecting mirrors and the beamsplitter has motivated dedicated search methods for dilaton dark matter (see Section III.1) and has been constrained using data from GEO600 and LIGO (see Section III.2). A search strategy targeting the temporal modulation of the laser frequency is still under development Hall and Aggarwal (2022).
@@ II.3.2 Changes to neutron-star moment of inertia
Beyond laboratory-based detectors, dilaton-induced fluctuations in fundamental constants could also have observable consequences in astrophysical systems. For example, they could alter moment of inertia of neutron stars Kaplan et al. (2022), which would produce deviations in the arrival time of pulses, to be discussed in Section III.3. In order to quantify the extent to which the neutron star moment of inertia will change, we must know how protons and neutrons will change in response to dilatons. This motivates us to consider the following interactions in QCD of quarks in the presence of dilatons Damour and Donoghue (2010); Kaplan et al. (2022):
@@
The shift in nucleon masses can then be obtained by noting that nucleons are composite objects whose masses arise mostly from QCD binding energy, with a smaller contribution from the quark masses. The dilaton couples to both the quark masses and the gluon field strength, so the fractional change in a nucleon mass is a weighted sum of these effects Damour and Donoghue (2010); Kaplan et al. (2022):
@@
where $I_{0}$ is the unperturbed moment of inertia and $\omega_{0}$ is the unperturbed angular frequency of the neutron star. Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro and Teukolsky (1983):
@@
The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons $N_{p}$ and neutrons $N_{n}$ (which dominate the pulsar’s mass). For neutron stars observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell et al. (2019); Kaplan et al. (2022).
@@
Using the above formalism, dilatons have been constrained with recent pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be extensively discussed in Section III.3.
@@ II.3.3 Changes to atomic clock times
Changes in fundamental constants would also affect the way in which atomic clocks measure time. Since pulsar timing array experiments reference their measurements of the ToAs of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum et al. (2004). Likewise, a shift in the frequency of the atomic clock will give rise to an apparent shift in the measured pulsar spin frequency, given by Eq. 38.
From Flambaum et al. (2004); Kaplan et al. (2022), the frequency of an atomic clock is:
@@
The fluctuations in fundamental constants will cause modulations in the frequencies of atomic clocks composed of atoms $A$. Using previously derived expressions for variations in the electron mass, fine-structure constant, and quark and nucleon masses, the fractional frequency shift can be obtained Flambaum et al. (2004); Kaplan et al. (2022):
@@
where $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$. For Cesium, $K_{A}=0.83$ depends on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum et al. (2004). The index $q$ runs over up and down quarks.
@@
where $\xi_{A}\equiv 4+K_{A}$, and $\hat{C}_{A}=C_{A}(m_{u}+m_{d})^{2}/2m_{u}m_{d}$. This type of DM interaction would leave imprints on pulsar timing array experiments Porayko et al. (2018); Smarra et al. (2023); Afzal et al. (2023), which will be discussed in Section III.3.
@@ II.4 Vector dark matter
So far, we have considered explain how spin-0 particles would affect GW interferometers; however, DM can also have nonzero spins. Here, we consider a model for DM that has spin-1, i.e. vectors, which we refer to as “dark photons”. Like dilatons, dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism Nelson and Scholtz (2011); Arias et al. (2012); Graham et al. (2016), parametric resonance or the tachyonic instability of a scalar field Agrawal et al. (2020); Co et al. (2019); Bastero-Gil et al. (2019); Dror et al. (2019), or from cosmic string network decays Long and Wang (2019). They could couple to standard-model particles, either to baryon or baryon-lepton number. In the interferometers, these interactions would occur everywhere, but would be most pronounced in the four interferometer test masses in the main resonant cavities of the instrument. The dark photons [^4] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen et al. (2023a). would cause a “dark” force on the mirrors, causing quasi-sinusoidal oscillations Pierce et al. (2018); Miller et al. (2021a), as will be shown below.
@@
which implies that we need to only consider the electric field. This dark electric field causes the test masses to oscillate at a quasi-fixed frequency given by the dark photon mass. The acceleration of a given mirror can thus be derived using Eqs. 49 and 48 Guo et al. (2019); Pierce et al. (2018):
@@
Assuming that all mirrors have the same charge-to-mass ratio, integrating LABEL:eqn:accel twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal can be computed Pierce et al. (2018):
@@
where $C=\sqrt{2}/3$ (for an interferometer with two perpendicular arms) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometries (the calculation for $C$ is shown in the appendix of Pierce et al. (2018) and in Morisaki et al. (2021)).
A second strain results due to the so-called “finite light travel time effect, in which the mirrors in the interferometers have moved in the time that it takes the light to reach them from the beamsplitter. This strain can actually be larger than that given in Eq. 54 and can be derived Morisaki et al. (2021):
@@ II.5 Tensor dark matter
Modifications to gravity due to an additional spin-2 particle could act as DM. Specifically, bimetric gravity Hassan and Rosen (2012); Schmidt-May and von Strauss (2016), a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola et al. (2018).
Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Fierz and Pauli (1939); Armaleo et al. (2020):
@@
where $M:=g^{\mu\nu}M_{\mu\nu}$ and $\mathcal{E}^{\mu\nu\rho\sigma}$ is the Lichnerowicz operator, defined in equation 2.2 of Armaleo et al. (2020).
For the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, the equations of motion for this field, assuming it is ultralight, can be derived at late times as in  Marzola et al. (2018); Aoki and Maeda (2018); Kun et al. (2018). The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position:
@@
The strain on GW detectors arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha_{\rm TB}$ [^5] In the literature, $\alpha_{\rm TB}=\alpha$ simply, but to disentangle it from the fine structure constants, we rename it. TB is tensor boson. Armaleo et al. (2020); thus, the perturbation $h_{ij}$ is:
@@
To obtain the strain on GW interferometers, we note that the detector response function can be parameterized in terms of a tensor, often written as $D^{ij}$ Romano and Cornish (2017). By contracting $h_{ij}$ with $D^{ij}$, the resulting strain can be derived Armaleo et al. (2021).
@@
We note that an interesting discrepancy arises when comparing the results from Refs. Aoki and Mukohyama (2016); Manita et al. (2024) with those from Refs. Marzola et al. (2018); Armaleo et al. (2021) regarding the strain induced by tensor dark matter. The two sets of works differ cause a factor of 2 difference in the strain, which can be traced back to a factor of eight difference in their respective Lagrangians. This difference stems from variations in the normalization factors and scaling of the tensor field amplitude, as well as differences in how the tensor field’s energy density is treated. Such discrepancies are not uncommon in the study of dark matter and may have implications for the interpretation of experimental constraints and the modeling of astrophysical signals. However, we note that no constraints on tensor DM have been placed as of yet using GW interferometers.

### v5

@@ II.2 Axions
Axions were originally proposed as a solution to the strong CP problem in QCD Peccei:1977hh; Peccei:1977ur, and more generally, axion-like particles arise in many extensions of the standard model. If sufficiently light and weakly interacting, they can also constitute a viable component of DM. A key feature of axions is their ability to couple to standard-model fields, in particular photons, through a Chern–Simons interaction. This coupling is described by the Lagrangian Chern:1974ft:
@@
the equations of motion for $A_{\text{L/R}}(t,\bm{k})$ can be derived Nagano:2019rbw:
@@
Eq. 11 implies that in the presence of an axion field, the phase velocities of circularly polarized photons in interferometers are modulated. In order to detect this effect in LIGO, Virgo and KAGRA instruments, additional but simplistic optical components are required in order to measure the optical path difference between two orthogonal polarizations of light ($p$- and $s$-polarizations) Nagano:2021kwx. In practice, GW interferometers use linearly polarized light ($p$-polarized), which becomes partially $s$-polarized due to axion-induced modulation. Similar deviations in polarization can also be probed with pulsar timing arrays Xue:2024zjq; EPTA:2024gxu, and experimental constraints on axions from them will be discussed in Section III.3.5. Notably, ground-based interferometers and pulsar timing arrays are sensitive to very different axion mass ranges, roughly $10^{-13},10^{-11}]$ eV and $10^{-23},10^{-21}]$ eV, respectively, but can achieve comparable absolute constraints on the axion-photon coupling $g_{\text{a}\gamma}$, which vary strongly with the axion mass. However, no constraints exist yet for this effect from ground-based interferometers.
@@ II.3 Dilaton scalar dark matter
Another class of ultralight dark matter candidates are dilatons, scalar fields that can couple to standard-model particles and induce time-dependent variations in fundamental constants Preskill:1982cy; Abbott:1982af; Dine:1982ah; Cho:1998aa; Cho:2007cy; Arvanitaki:2014faa; Stadnik2015a; Stadnik2015b; Stadnik2016. In particular, a dilaton field can cause the electron mass and other physical constants to oscillate at the dilaton frequency. Similarly to the axion, the scalar ultralight dark-matter field $\phi$ can be written as Arvanitaki:2014faa; Derevianko:2016vpm; Vermeulen:2021epa:
@@
where $\vec{x}$ is a position vector, $\vec{k}=m_{\rm DM}\vec{v}_{0}$ is the wave vector, and $\Psi$ is a phase factor, and has a corresponding Lagrangian $\mathcal{L}_{\mathrm{int}}$ Vermeulen:2021epa:
@@ II.3.1 Changes to size and indices of refraction of solids
The oscillations in fundamental constants induced by a dilaton field lead to modulations of the Bohr radii of atoms in interferometer components grote2019novel, leading to measurable changes in the size and refractive index of the beam splitter Vermeulen:2021epa, and the arm mirrors Gottel:2024cfj. Oscillations in the beam splitter size cause photons traveling down each arm to traverse slightly different paths along its surface, producing a differential phase shift. Initially, it was assumed that the beam-splitter effect was the dominant contribution Vermeulen:2021epa, but more recent work has shown that small differences in the thicknesses of the mirrors in each arm also lead to non-negligible path-length variations in the interferometer Gottel:2024cfj.
@@
In the context of GW interferometers, this effect translates into a differential displacement of the beam splitter along the two arms. For the $x$-arm, assuming light incidence at 45 degrees and 50% mirror reflectivity, the induced change in arm length, $\delta L_{x}$, can be computed as in grote2019novel:
@@
where $w$ is the thickness of the mirrors, $n$ is the index of the beamsplitter, and $l$ is the length of the beamsplitter. Along the $y$-arm, the change in arm length is Vermeulen:2021epa:
@@
We have only considered the contribution arising from the change in length of the beamsplitter, not from the change in index of refraction, because the former is much larger than the latter grote2019novel.
In addition to dilatons interacting with atoms in the beamsplitter, the reference cavity of the detector can provide a region in which dilatons interact appreciably with light, allowing for potential detection. Laser light locked to an optical cavity made out of a solid material will have its frequency modulated due to the presence of a dilaton DM field, compared to light locked to a free-space suspended cavity, whose frequency would remain stable Derevianko:2016vpm; Geraci:2018fax; Hall:2022zvi. The strain amplitude of such a signal, which would be distinct from that in Eq. 26, can be written as:
@@
The effects described in Eq. 26 and Eq. 27 can both produce measurable signals in interferometers. Currently, however, only the spatial strain affecting mirrors and the beamsplitter has motivated dedicated search methods for dilaton dark matter (see Section III.1) and has been constrained using data from GEO600 and LIGO (see Section III.2). A search strategy targeting the temporal modulation of the laser frequency is still under development Hall:2022zvi.
@@ II.3.2 Changes to neutron-star moment of inertia
Beyond laboratory-based detectors, dilaton-induced fluctuations in fundamental constants could also have observable consequences in astrophysical systems. For example, they could alter moment of inertia of neutron stars Kaplan:2022lmz, which would produce deviations in the arrival time of pulses, to be discussed in Section III.3. In order to quantify the extent to which the neutron star moment of inertia will change, we must know how protons and neutrons will change in response to dilatons. This motivates us to consider the following interactions in QCD of quarks in the presence of dilatons Damour:2010rp; Kaplan:2022lmz:
@@
The shift in nucleon masses can then be obtained by noting that nucleons are composite objects whose masses arise mostly from QCD binding energy, with a smaller contribution from the quark masses. The dilaton couples to both the quark masses and the gluon field strength, so the fractional change in a nucleon mass is a weighted sum of these effects Damour:2010rp; Kaplan:2022lmz:
@@
where $I_{0}$ is the unperturbed moment of inertia and $\omega_{0}$ is the unperturbed angular frequency of the neutron star. Assuming a simplistic model of a non-rotating, spherically symmetric neutron star composed only of neutrons with a polytropic equation-of-state, the radius and moment of inertia can be written as Shapiro:1983du:
@@
The sum is over the different types of particles present in the pulsar: electrons, muons, protons and neutrons, $Y_{f}\equiv N_{f}/(N_{n}+N_{p})$ is a weighting function that gives the number of $f$ particles relative to the sum of protons $N_{p}$ and neutrons $N_{n}$ (which dominate the pulsar’s mass). For neutron stars observed by pulsar timing arrays, the following is assumed for $Y_{f}$: $Y_{n}=0.9$, $Y_{p}=Y_{e}=0.1$ and $Y_{\mu}=0.05$ Bell:2019pyc; Kaplan:2022lmz.
@@
Using the above formalism, dilatons have been constrained with recent pulsar timing array experiments Porayko:2018sfa; EuropeanPulsarTimingArray:2023egv; NANOGrav:2023hvm, which will be extensively discussed in Section III.3.
@@ II.3.3 Changes to atomic clock times
Changes in fundamental constants would also affect the way in which atomic clocks measure time. Since pulsar timing array experiments reference their measurements of the ToAs of pulses to terrestrial time (TT) using atomic clocks composed primarily of cesium, any oscillation of the fundamental constants will affect the atomic clocks, and thus the times measured Flambaum:2004tm. Likewise, a shift in the frequency of the atomic clock will give rise to an apparent shift in the measured pulsar spin frequency, given by Eq. 38.
From Flambaum:2004tm; Kaplan:2022lmz, the frequency of an atomic clock is:
@@
The fluctuations in fundamental constants will cause modulations in the frequencies of atomic clocks composed of atoms $A$. Using previously derived expressions for variations in the electron mass, fine-structure constant, and quark and nucleon masses, the fractional frequency shift can be obtained Flambaum:2004tm; Kaplan:2022lmz:
@@
where $\delta F_{\rm rel}/F_{\rm rel}=K_{A}$, and the derivative of the nuclear magnetic moment is $\delta\mu/\mu=C_{A}\,\delta m_{q}/m_{q}$. For Cesium, $K_{A}=0.83$ depends on $Z$ and $\alpha$, and $C_{A}=0.110$ Flambaum:2004tm. The index $q$ runs over up and down quarks.
@@
where $\xi_{A}\equiv 4+K_{A}$, and $\hat{C}_{A}=C_{A}(m_{u}+m_{d})^{2}/2m_{u}m_{d}$. This type of DM interaction would leave imprints on pulsar timing array experiments Porayko:2018sfa; EuropeanPulsarTimingArray:2023egv; NANOGrav:2023hvm, which will be discussed in Section III.3.
@@ II.4 Vector dark matter
So far, we have considered explain how spin-0 particles would affect GW interferometers; however, DM can also have nonzero spins. Here, we consider a model for DM that has spin-1, i.e. vectors, which we refer to as “dark photons”. Like dilatons, dark photons could explain the entirety of the relic abundance of DM, which could arise from the misalignment mechanism nelson2011dark; arias2012wispy; graham2016vector, parametric resonance or the tachyonic instability of a scalar field agrawal2020relic; Co:2018lka; Bastero-Gil:2018uel; Dror:2018pdh, or from cosmic string network decays long2019dark. They could couple to standard-model particles, either to baryon or baryon-lepton number. In the interferometers, these interactions would occur everywhere, but would be most pronounced in the four interferometer test masses in the main resonant cavities of the instrument. The dark photons [^4] These are different dark photons than those that kinetically mix with the ordinary photon, which will be discussed later in the context of luminous superradiance Siemonsen:2022ivj. would cause a “dark” force on the mirrors, causing quasi-sinusoidal oscillations Pierce:2018xmy; Miller:2020vsl, as will be shown below.
@@
which implies that we need to only consider the electric field. This dark electric field causes the test masses to oscillate at a quasi-fixed frequency given by the dark photon mass. The acceleration of a given mirror can thus be derived using Eqs. 49 and 48 Guo:2019ker; Pierce:2018xmy:
@@
Assuming that all mirrors have the same charge-to-mass ratio, integrating LABEL:eqn:accel twice over time, and averaging over random polarization and propagation directions, the strain on the interferometers caused by a dark photon DM signal can be computed Pierce:2018xmy:
@@
where $C=\sqrt{2}/3$ (for an interferometer with two perpendicular arms) is a geometrical factor obtained by averaging over all possible dark photon propagation and polarization directions, and different detector geometries (the calculation for $C$ is shown in the appendix of Pierce:2018xmy and in Morisaki:2020gui).
A second strain results due to the so-called “finite light travel time effect, in which the mirrors in the interferometers have moved in the time that it takes the light to reach them from the beamsplitter. This strain can actually be larger than that given in Eq. 54 and can be derived Morisaki:2020gui:
@@ II.5 Tensor dark matter
Modifications to gravity due to an additional spin-2 particle could act as DM. Specifically, bimetric gravity Hassan:2011zd; Schmidt-May:2015vnx, a theory in which a massless and massive spin-2 field interact, provides a plausible candidate for ultralight dark matter Marzola:2017lbt.
Here, we consider a massive spin-2 field $M_{\mu\nu}$ described by the Fierz-Pauli Lagrangian density Fierz:1939ix; Armaleo:2020yml:
@@
where $M:=g^{\mu\nu}M_{\mu\nu}$ and $\mathcal{E}^{\mu\nu\rho\sigma}$ is the Lichnerowicz operator, defined in equation 2.2 of Armaleo:2020yml.
For the Friedman-Lemaitre-Robertson-Walker (FLRW) background metric, the equations of motion for this field, assuming it is ultralight, can be derived at late times as in  Marzola:2017lbt; Aoki:2017cnz; Kun:2018ino. The field spatial component can be written in a similar way as for dark photons and scalar bosons at given position:
@@
The strain on GW detectors arises analogously to that from GWs: a stretching of space-time in the presence of the field, since the massive spin-2 metric and its coupling constant can be absorbed into the definition of the massless spin-2 metric in the linear regime of the coupling constant $\alpha_{\rm TB}$ [^5] In the literature, $\alpha_{\rm TB}=\alpha$ simply, but to disentangle it from the fine structure constants, we rename it. TB is tensor boson. Armaleo:2020yml; thus, the perturbation $h_{ij}$ is:
@@
To obtain the strain on GW interferometers, we note that the detector response function can be parameterized in terms of a tensor, often written as $D^{ij}$ Romano:2016dpx. By contracting $h_{ij}$ with $D^{ij}$, the resulting strain can be derived Armaleo:2020efr.
@@
We note that an interesting discrepancy arises when comparing the results from Refs. Aoki:2016zgp; Manita:2023mnc with those from Refs. Marzola:2017lbt; Armaleo:2020efr regarding the strain induced by tensor dark matter. The two sets of works differ cause a factor of 2 difference in the strain, which can be traced back to a factor of eight difference in their respective Lagrangians. This difference stems from variations in the normalization factors and scaling of the tensor field amplitude, as well as differences in how the tensor field’s energy density is treated. Such discrepancies are not uncommon in the study of dark matter and may have implications for the interpretation of experimental constraints and the modeling of astrophysical signals. However, we note that no constraints on tensor DM have been placed as of yet using GW interferometers.